# Measuring Prosody Diversity in Zero-Shot TTS: A New Metric, Benchmark, and Exploration

### v2

@@ Abstract
Index Terms—  Zero-shot TTS, prosody diversity, prosody evaluation, prosody benchmark
@@ 1 Introduction
Prosody plays a central role in spoken communication, conveying paralinguistic information such as emotion, attitude, and intent that shapes how listeners interpret meaning. Subtle variations in pitch, intensity, and duration can fundamentally alter the interpretation of an utterance even with identical text, making prosody crucial to naturalness and expressiveness in speech synthesis. While zero-shot TTS [1, 2, 3] has advanced in intelligibility and speaker similarity, prosody diversity has received far less attention. Existing evaluation practices [4] rely largely on log $F_{0}$ root mean squared error (RMSE), which correlates weakly with human judgments, captures pitch but neglects rhythm and intensity [5], and requires costly dynamic time warping (DTW). Consequently, a clear gap remains between automatic metrics and human perception of prosody diversity.
@@
Recent efforts have shown that both continuous [6] and discrete [7] self-supervised (SSL) speech representations effectively encode prosodic information. Concurrently, NLP automatic evaluation metrics have been adapted to discrete speech representations for reference-free [8] and reference-aware [9] speech quality assessment. In particular, SpeechTokenDistance [9] is a simple approach but remains preliminary, as it focuses solely on measuring correlation with acoustic metrics without examining what the differences represent or further analyzing correlations with human perception.
@@
With this perspective in mind, we introduce ProsodyEval, the first human-annotated dataset for prosody diversity assessment, and DS-WED (Discretized Speech Weighted Edit Distance), a new objective diversity metric that quantifies prosodic variation via weighted edit distance over semantic token sequences derived from speech SSL models such as HuBERT [10] and WavLM [11] through $k$-kmeans clustering. ProsodyEval contains 1000 synthetic samples from 7 mainstream TTS systems, paired with 2000 human ratings of prosody diversity. Experiments show that DS-WED achieves substantially higher correlation with human judgments than frequently used acoustic metrics, including log $F_{0}$ RMSE and Mel cepstral distortion (MCD), while remaining robust across models, layers, and cluster sizes. Building on DS-WED, we establish the first systematic benchmark of prosody diversity across state-of-the-art open-source TTS systems on LibriSpeech test-clean and Seed-TTS test-en. Our further explorations reveal the impact of modeling paradigms, duration control, and reinforcement learning (RL), and demonstrate that even advanced large audio language models (LALMs) like Gemini 2.5 Pro remain unreliable for prosody evaluation.
Our contributions to the community include four aspects:
  • Dataset. We present ProsodyEval, the first human-annotated dataset for prosody diversity assessment in zero-shot TTS.
@@
  • Metrics. We propose DS-WED, a new objective prosody diversity metric based on semantic token weighted edit distance that better correlates with human ratings than existing acoustic metrics.
@@
  • Benchmark. We establish the first benchmark of prosody diversity across state-of-the-art open-source TTS systems, offering systematic comparisons and analyses of different generative paradigms from the perspective of prosody diversity.
@@
  • Exploration. We investigate key factors shaping prosody diversity, showing that (1) autoregressive (AR) systems outperform flow-matching based non-autoregressive (NAR) systems but not masked generative modeling (MGM); (2) duration variation is critical, yet NAR systems lack explicit duration control, and flow-matching models with implicit alignment suffer from inherent constraints; (3) RL via direct preference optimization (DPO) trades off diversity for intelligibility; and (4) current LALMs remain unreliable for prosody understanding.
  • All related resources, including data, toolkit, benchmark, and $k$-means models, will be released to facilitate future research.
@@ 2.1 Data Collection
We construct the ProsodyEval dataset by aggregating synthetic speech from diverse generative paradigms, spanning AR and NAR approaches, including next-token prediction [12, 13, 3], flow matching [14, 15, 16], and MGM [17]. Concretely, ProsodyEval comprises samples synthesized by seven recent open-source TTS systems, namely XTTS-v2 [12], CosyVoice [13], CosyVoice 2 [13], E2 TTS [14], F5-TTS [15], MaskGCT [17], and ZipVoice [16]. These paradigms reflect widely used approaches in modern TTS, rendering the dataset representative for evaluating prosodic variation.
@@ 2.3 Acoustic Prosody Metrics
Table 2: Prosody diversity benchmark of zero-shot TTS systems from diverse generative paradigms on LibriSpeech test-clean and Seed-TTS test-en, using traditional acoustic metrics log $F_{0}$ RMSE and MCD as well as our proposed DS-WED, each computed as micro-average (Avg.) and rank-based score (Borda Avg.). The best results are highlighted in bold, and second-best are underlined.
@@ 3 DS-WED Metric
Formulation   Let $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ denote two synthesized speech samples. To remove the influence of leading and trailing silences, a pretrained VAD model [21] is employed to trim the raw waveforms:
@@
$\mathbf{c}_{1}=\mathrm{Encode}_{\mathrm{spch}}(\tilde{\mathbf{X}}_{1}),\qquad\mathbf{c}_{2}=\mathrm{Encode}_{\mathrm{spch}}(\tilde{\mathbf{X}}_{2}).$ (2)
@@
$\mathrm{\text{DS-WED}}(\mathbf{c}_{1},\mathbf{c}_{2})=\min_{\pi\in\mathcal{A}(\mathbf{c}_{1},\mathbf{c}_{2})}\sum_{(i,j,o)\in\pi}w_{o}\,c_{o}(c_{1,i},c_{2,j}),$ (3)
where $\pi$ is an edit path in the alignment set $\mathcal{A}$, $o\in\{\mathrm{sub},\mathrm{ins},\mathrm{del}\}$ is the edit operation, $c_{o}(\cdot)$ denotes the corresponding edit cost, and $w_{o}$ represents an operation-dependent weight.
@@
  • Operation weights can be tuned to perceptual sensitivity. Based on listening tests showing more sensitivity to intonation and word stress than to pause duration, we increase $w_{\mathrm{sub}}$ to 1.2 while keeping $w_{\mathrm{ins}}$ and $w_{\mathrm{del}}$ at 1. The weighted edit distance can be interpreted as the minimum perceptually prosodic modifications needed to transform one speech into another at the discrete level.
@@ 5 Conclusion
In this work, we introduce ProsodyEval, the first human-annotated dataset for assessing prosody diversity in zero-shot TTS, together with DS-WED, a new objective diversity metric based on weighted edit distance over semantic tokens. DS-WED outperforms frequently used acoustic metrics in correlating with human judgments and remains robust in speech discretization across models, layers, and cluster sizes, enabling more reliable evaluation of prosodic variation. Leveraging DS-WED, we systematically benchmark and analyze mainstream zero-shot TTS systems. Our findings indicate that (1) AR systems outperform NAR flow-matching models but not MGM, (2) duration variation strongly shapes prosody diversity, yet NAR systems lack duration control, and flow-matching models with implicit alignment suffer from inherent architectural limitations, (3) RL via DPO improves intelligibility while reducing prosody diversity, and (4) even advanced LALMs like Gemini 2.5 Pro remain unreliable for prosody evaluation. One limitation lies in the cross-lingual applicability of DS-WED, validated only on English. Looking ahead, ProsodyEval and DS-WED fill a gap in zero-shot TTS evaluation and open new avenues for speech data engineering.

### v3

@@ Abstract
Index Terms—  Zero-shot TTS, prosody diversity, prosody diversity evaluation, prosody diversity benchmark
@@ 1 Introduction
Prosody plays a central role in spoken communication, conveying paralinguistic information such as emotion, attitude, and intent that shapes how listeners interpret meaning. Subtle variations in pitch, intensity, and duration can fundamentally alter the interpretation of an utterance even with identical text, making prosody crucial to naturalness and expressiveness in speech synthesis. While zero-shot TTS [8, 25, 29, 26] has advanced in intelligibility and speaker similarity, prosody diversity has received far less attention. Existing evaluation practices [13] rely largely on log $F_{0}$ root mean squared error (RMSE), which correlates weakly with human judgments, captures pitch but neglects rhythm and intensity [2], and requires costly dynamic time warping (DTW). Consequently, a clear gap remains between automatic metrics and human perception of prosody diversity.
@@
Recent efforts have shown that both continuous [19] and discrete [21] self-supervised (SSL) speech representations [31] effectively encode prosodic information. Concurrently, NLP automatic evaluation metrics have been adapted to discrete speech representations for reference-free [20] and reference-aware [23] speech quality assessment. In particular, SpeechTokenDistance [23] is a simple approach but remains preliminary, as it focuses solely on measuring correlation with acoustic metrics without examining what the differences represent or further analyzing correlations with human perception.
@@
With these perspectives in mind, we introduce ProsodyEval, a human-annotated dataset for prosody diversity assessment, and DS-WED (Discretized Speech Weighted Edit Distance), a new objective diversity metric that quantifies prosodic variation via weighted edit distance over semantic token sequences derived from speech SSL models such as HuBERT [16] and WavLM [4] through $k$-kmeans clustering. ProsodyEval contains 1000 synthetic samples from 7 mainstream TTS systems, paired with 2000 human ratings of prosody diversity. Experiments show that DS-WED achieves substantially higher correlation with human judgments than frequently used acoustic metrics, including log $F_{0}$ RMSE and Mel cepstral distortion (MCD), while remaining robust across models, layers, and cluster sizes. Building on DS-WED, we establish the first systematic benchmark of prosody diversity across state-of-the-art open-source TTS systems on LibriSpeech test-clean and Seed-TTS test-en. Our further explorations reveal the impact of modeling paradigms, duration control, and reinforcement learning (RL), and demonstrate that even advanced large audio language models (LALMs) like Gemini 2.5 Pro remain unreliable for prosody evaluation.
Our contributions [^1] Code and models are publicly available at yfyeung/DS-WED. are threefold:
@@
  • We propose DS-WED, a reliable objective prosody diversity metric based on semantic token weighted edit distance. Through validation on ProsodyEval, a human-annotated dataset for prosody diversity in zero-shot TTS, we demonstrate DS-WED correlates better with human ratings than existing acoustic metrics.
@@
  • We establish the first benchmark of prosody diversity across state-of-the-art open-source TTS systems, offering systematic comparisons and analyses of different generative paradigms.
@@
  • We investigate key factors shaping prosody diversity, showing that (1) autoregressive (AR) systems outperform flow-matching based non-autoregressive (NAR), but not masked generative modeling (MGM), as duration variation is critical; however, NAR lacks duration modeling, while flow-matching models with implicit alignment suffer from inherent constraints; (2) RL via direct preference optimization (DPO) trades off diversity for intelligibility; and (3) current LALMs remain unreliable for prosody understanding.
@@ 2.1 Data Collection
We construct ProsodyEval dataset by aggregating synthetic speech from diverse generative paradigms, spanning AR and NAR approaches, including next-token prediction [3, 9, 11], flow matching [12, 5, 33], and MGM [27]. Concretely, ProsodyEval comprises samples synthesized by seven recent open-source TTS systems, namely XTTS-v2 [3], CosyVoice [9], CosyVoice 2 [9], E2 TTS [12], F5-TTS [5], MaskGCT [27], and ZipVoice [33]. These paradigms reflect widely used approaches in modern TTS, rendering the dataset representative for evaluating prosodic variation.
@@ 2.3 Acoustic Prosody Metrics
Table 2: Prosody diversity benchmark of zero-shot TTS systems from diverse generative paradigms on LibriSpeech test-clean and Seed-TTS test-en, using traditional acoustic metrics log $F_{0}$ RMSE and MCD as well as our proposed DS-WED, each computed as micro-average (Avg.) and rank-based score (Borda Avg.). The best results are highlighted in bold, and the second-best are underlined.
@@ 3 DS-WED Metric
Formulation   Let $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ denote two synthesized speech samples. To remove the influence of leading and trailing silences, a pre-trained VAD model [^2] https://github.com/snakers4/silero-vad is employed to trim the raw waveforms:
@@
$\mathbf{c}_{1}=\mathrm{Encode}_{\mathrm{spch}}(\tilde{\mathbf{X}}_{1}),\qquad\mathbf{c}_{2}=\mathrm{Encode}_{\mathrm{spch}}(\tilde{\mathbf{X}}_{2}).\vskip-2.77501pt$ (2)
@@
$\mathrm{\text{DS-WED}}(\mathbf{c}_{1},\mathbf{c}_{2})=\min_{\pi\in\mathcal{A}(\mathbf{c}_{1},\mathbf{c}_{2})}\sum_{(i,j,o)\in\pi}w_{o}\,c_{o}(c_{1,i},c_{2,j}),\vskip-2.77501pt$ (3)
where $\pi$ is an edit path in the alignment set $\mathcal{A}$, $o\in\{\mathrm{sub},\mathrm{ins},\mathrm{del}\}$ is the edit operation, $c_{o}(\cdot)$ denotes the corresponding edit cost, and $w_{o}$ is an operation-dependent weight, set to $1$ in all experiments.
@@
  • Operation weights can be tuned to perceptual sensitivity. Listening tests indicate higher sensitivity to intonation and word stress than to pause duration, suggesting that assigning larger weights to substitution operations may better align with human judgment.
  • The weighted edit distance can be interpreted as the minimum perceptually prosodic modifications needed to transform one speech into another at the discrete level.
@@ 5 Conclusion
In this work, we propose DS-WED, a reliable objective diversity metric based on weighted edit distance over semantic tokens, together with ProsodyEval, a human-annotated dataset for assessing prosody diversity in zero-shot TTS. DS-WED outperforms frequently used acoustic metrics in correlating with human judgments and remains robust in speech discretization across models, layers, and cluster sizes, enabling reliable quantitative evaluation of prosodic variation. Leveraging DS-WED, we systematically benchmark and analyze mainstream zero-shot TTS systems. We quantitatively show that (1) AR systems outperform NAR flow-matching models but not MGM, as duration variation strongly shapes prosody diversity; however, NAR lack duration modeling, and flow-matching models with implicit alignment suffer from inherent limitations, (2) RL via DPO improves intelligibility while reducing prosody diversity, and (3) even advanced LALMs like Gemini 2.5 Pro remain unreliable for prosody evaluation. One limitation lies in the cross-lingual applicability of DS-WED, validated only on English. Looking ahead, ProsodyEval and DS-WED fill a gap in zero-shot TTS evaluation and open new avenues for speech data engineering.